# Simulated host galaxy analogs of high-z quasars observed with JWST

### v1

@@ 7 Conclusions
With the criteria described in section 4.1, we select two analogs similar to J2255+0251 or J2236+0032 in stellar mass, black hole mass, and galactic radius. We compare the BlueTides analogs to the actual JWST detected quasar hosts and find the photometric galaxy magnitudes match closely between simulation and observation. We use mock imaging and a point source subtraction algorithm to make qualitative comparisons and begin to understand the effects of observational biases with these two analogs. We find that the mock images of BlueTides analogs are qualitatively similar to the detected hosts in Ding et al. (2023). However, the host detection remains difficult. This supports the lack of quasar host detections in Stone et al. (2023a). We do detect all hosts, but note that the success is highly dependent on the luminosity of the host galaxy. When we determine the half light radius and magnitudes for mock images of galaxies, we find that the recovered magnitudes agree mostly to within one sigma. The radii have a larger discrepancy between actual and post point source removal values with a maximum percent difference of or about 0.4 pkpc. The point source removal process consistently underestimates the radii in the galaxies tested in Table 4. This motivates future work to quantify these biases.
@@ Acknowledgements
We are grateful to Xuheng Ding and the rest of the SHELLQs team for helpful email communication and allowing the use of their JWST images in this paper. We also thank Mira Mechtley who wrote the point source removal software, psfmc. S.B. thanks Yuxiang Qin and Xuejian Shen for helpful comments or code that contributed to this paper.

### v3

@@ 7 Conclusions
With the criteria described in section 4.1, we select two analogs similar to J2255+0251 or J2236+0032 in stellar mass, black hole mass, and galactic radius. We compare the BlueTides analogs to the actual JWST detected quasar hosts and find the photometric galaxy magnitudes match closely between simulation and observation. We use mock imaging and a point source subtraction algorithm to make qualitative comparisons and begin to understand the effects of observational biases with these two analogs. We find that the mock images of BlueTides analogs are qualitatively similar to the detected hosts in Ding et al. (2023). However, the host detection remains difficult. This supports the lack of quasar host detections in Stone et al. (2024). We do detect all hosts, but note that the success is highly dependent on the luminosity of the host galaxy. When we determine the half light radius and magnitudes for mock images of galaxies, we find that the recovered magnitudes agree mostly to within one sigma. The radii have a larger discrepancy between actual and post point source removal values with a maximum percent difference of or about 0.4 pkpc. The point source removal process consistently underestimates the radii in the galaxies tested in Table 4. This motivates future work to quantify these biases.
@@
Data Availability
Data from the BlueTides simulation is available at http://bluetides.psc.edu. Other data generated in this work will be shared on reasonable request to the corresponding author. The reproduced Ding et al. (2023) images in this article were provided by Springer Nature with permission.
@@ Acknowledgements
We thank the anonymous referee for their comments which greatly helped to improve this work.
We are grateful to Xuheng Ding and the rest of the SHELLQs team for helpful email communication and allowing the use of their JWST images in this paper. We also thank Mira Mechtley who wrote the point source removal software, psfmc. S.B. thanks Chris Lovell, Yuxiang Qin and Xuejian Shen for helpful comments or code that contributed to this paper.